OpenStudy (ttop0816):

Solve for c. ab + c = 5 c = 5/ab <--- right?? c = 5 - ab c = 5 + ab

OpenStudy (anonymous):

well, first you divide both sides by ab, since youre solving for c, so yes youre Correct :)

OpenStudy (anonymous):

not correct ^

OpenStudy (ttop0816):

thank you! do you mind helping with another question?

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

you subtract im so sorry!!

OpenStudy (anonymous):

You must subtract ab from the left side.

OpenStudy (anonymous):

c = 5-ab

OpenStudy (anonymous):

^^ yeah, im stupid-

OpenStudy (ttop0816):

oh.. then it will be c = 5 - ab?

OpenStudy (anonymous):

yes

OpenStudy (ttop0816):

Solve for b. 3a + 2b = c b = c - (3a - 2) b = (c - 3a)/2 <--- would it be this?? b = (3a - c)/2

OpenStudy (anonymous):

first you subtract 3a from both sides, then divide both sides by 2

OpenStudy (anonymous):

so yes, youre correct!